# The OpenHands Software Agent SDK: A Composable and Extensible Foundation for Production Agents

### v1

@@ 6 Related Work
Recent agent SDKs share common goals of tool orchestration, state management, and production deployment. General-purpose frameworks such as LangChain and LangGraph focus on compositional pipelines and stateful graph execution with durable checkpoints for long-running reasoning workflows (LangChain, 2025; Team, 2025). Provider SDKs from OpenAI and Anthropic emphasize production orchestration, guardrails, and handoffs within their own model ecosystems (OpenAI, 2025; Anthropic, 2025a), while Google’s Agent Development Kit (ADK) targets model-agnostic agents integrated with Vertex AI for managed deployment (Google, 2025).
In contrast, the OpenHands Software Agent SDK is a fully open-source, vendor-agnostic platform purpose-built for software engineering agents. It couples an event-sourced state model and immutable configuration with sandboxed execution, a built-in REST/WebSocket server, and workspace-level remote interfaces (VS Code, VNC, browser). These design choices enable reproducible, deterministic execution and seamless transition from local prototyping to production-scale deployments—capabilities not addressed by existing general-purpose SDKs. A feature-level comparison with major agent SDKs is provided in Tab. 3.
@@ 7 Conclusion
The OpenHands Software Agent SDK bridges the gap between rapid local prototyping and production deployment for software engineering agents through a stateless, event-sourced, and composable architecture spanning four packages (SDK, Tools, Workspace, Server). From our experience scaling OpenHands from research prototype to production system, we learned that strict separation between core agent logic and applications is essential for maintainability; event-sourced state enables reproducibility and fault recovery; immutable component design prevents configuration drift; and a unified execution model that supports both local development and sandboxed deployment streamlines the transition from experimentation to production. These lessons collectively shaped a system that supports reliable, extensible operation across heterogeneous environments. Evaluation on SWE-Bench Verified and GAIA benchmarks confirms strong performance and consistency across diverse model backends, validating the SDK as a robust foundation for both research and industrial-scale deployment.

### v2

@@ 6 Related Work
Recent agent SDKs share common goals of tool orchestration, state management, and production deployment. General-purpose frameworks such as LangChain and LangGraph focus on compositional pipelines and stateful graph execution with durable checkpoints for long-running reasoning workflows (langchain; langgraph). LangGraph Platform further offers deployment and server capabilities, making it one of the most feature-complete open-source alternatives. Provider SDKs from OpenAI and Anthropic emphasize production orchestration, guardrails, and handoffs within their own model ecosystems (openai-agents; anthropic-agents), while Google’s Agent Development Kit (ADK) targets model-agnostic agents integrated with Vertex AI for managed deployment (google-adk).
In contrast, the OpenHands Software Agent SDK is a fully open-source, vendor-agnostic platform purpose-built for software engineering agents. It couples an event-sourced state model and immutable configuration with sandboxed execution, a built-in REST/WebSocket server, and workspace-level remote interfaces for human-agent collaboration. These design choices enable reproducible, deterministic execution and seamless transition from local prototyping to production-scale deployments. A feature-level comparison with major agent SDKs—including LangChain/LangGraph—is provided in Tab. 6, with comparison methodology detailed in the caption.
7 Discussion
Architectural Alternatives Considered.
During the V1 redesign, we evaluated several alternative approaches. For state management, we considered a traditional database-backed model (used by many web frameworks) but rejected it because it would couple the SDK to a specific storage backend and make offline replay difficult; event sourcing was chosen for its reproducibility and storage-agnostic design. For execution isolation, we considered mandatory containerization (as in V0) and fully local-only execution; the optional isolation model was selected as a compromise that aligns with MCP’s local-first assumptions while preserving sandboxing for production. For the package structure, we considered a monolithic SDK (simpler dependency management) versus a microservices architecture (maximum flexibility); the four-package design balances independent testability with practical dependency management.
Multi-Tenant Deployment.
The current security and isolation discussion focuses primarily on single-user scenarios. In multi-tenant deployments—where multiple users submit requests to the same agentic application or multiple applications share infrastructure components—additional considerations apply. The containerized deployment model (§4.10) provides process-level isolation between tenants, with each agent instance running in an independent container with a dedicated filesystem and resource limits. The event-sourced state model naturally supports tenant isolation since each conversation maintains its own event log. However, shared resources such as LLM API keys, MCP servers, and secret registries require careful access control. The SDK’s SecretRegistry supports per-conversation secrets, and the agent server’s session-based authentication provides a foundation for multi-tenant access control, but a comprehensive multi-tenant security audit remains future work.
Limitations.
The current implementation focuses on single-agent conversations. While the event-sourced architecture naturally supports interleaving events from multiple agents, coordination mechanisms for multi-agent collaboration require further design. The security framework, while substantially improved over V0, cannot guarantee complete safety: LLM-based security analysis is subject to adversarial prompts and inconsistent classification.
@@ 8 Conclusion
The OpenHands Software Agent SDK provides a reference architecture for production agent systems, bridging the gap between rapid local prototyping and production deployment through a stateless, event-sourced, and composable design spanning four packages (SDK, Tools, Workspace, Server). From our experience scaling OpenHands from research prototype to production system, we distill four key lessons: strict separation between core agent logic and applications is essential for maintainability; event-sourced state enables reproducibility and fault recovery with negligible overhead (sub-millisecond persist latency, crash recovery under 20 ms); immutable component design prevents configuration drift; and a unified execution model supporting both local and sandboxed deployment streamlines the transition from experimentation to production. These architectural choices yield measurable reliability gains: a 15-day production comparison demonstrates a 61% reduction in system-attributable failures relative to V0, while evaluations across 14 models and five benchmark categories confirm that the redesign preserves and enables strong agentic performance, achieving state-of-the-art results on three of five tasks. Together, these results validate the SDK as a robust foundation for both research and industrial-scale deployment.